The natural history of religion^{∗}

by Winwood Reade (1838–1875)

1872

When the poet invokes in his splendid frenzy the shining spheres of heaven, the murmuring fountains, and the rushing streams; when he calls upon the earth to hearken, and bids the wild sea listen to his song; when he communes with the sweet secluded valleys and the haughty-headed hills, as if those inanimate objects were alive, as if those masses of brute matter were endowed with sense and thought, we do not smile, we do not sneer, we do not reason, but we feel. A secret chord is touched within us; a slumbering sympathy is awakened into life. Who has not felt an impulse of hatred, and perhaps expressed it in a senseless curse, against a fiery stroke of sunlight, or a sudden gust of wind? Who has not felt a pang of pity for a flower torn and trampled in the dust; a shell dashed to fragments by the waves? Such emotions or ideas last only for a moment; they do not belong to us; they are the fossil fancies of a bygone age; they are a heritage of thought from the childhood of our race. For there was a time when they possessed the human mind. There was a time when the phrases of modern poetry were the facts of ordinary life. There was a time when man lived in fellowship with nature, believing that all things which moved or changed had minds and bodies kindred to his own.

To those primeval people the sun was a great being, who brightened them in his pleasure, and who scorched them in his wrath. The earth was a sleeping monster; sometimes it rose a little and turned itself in bed. They walked upon its back when living; they were put into its belly when they died. Fire was a savage animal, which bit when it was touched. The birds and beasts were foreigners, possessing languages and customs of their own. The plants were dumb creatures, with characters good or bad, sometimes gloomy in aspect, malignant in their fruit, sometimes dispensing wholesome food and pleasant shade.

These various forms of nature they treated precisely as if they had been men. They sometimes adorned a handsome tree with bracelets like a girl; they offered up prayers to the fruit trees, and made them presents to coax them to a liberal return. They forbade the destruction of certain animals which they revered on account of their wisdom, or feared on account of their fierceness, or valued on account of their utility. They submitted to the tyranny of the more formidable beasts of prey, never venturing to attack them, for fear the nation of species should retaliate; but made them propitiatory gifts. In the same manner they offered sacrifices to avert the fury of the elements, or in gratitude for blessings which had been bestowed. But often a courageous people, when invaded, would go to war, not only with the tiger and the bear, but with powers which, to them, were not less human-like and real. They would cut with their swords at the hot wind of the desert, hurl their spears into the swollen river, stab the earth, flog the sea, shoot their arrows at the flashing clouds, and build up towers to carry heaven by assault.

But when through the operation of the Law of Growth the intellectual faculties of men become improved, they begin to observe their own nature, and in course of time a curious discovery is made. They ascertain that there is something which resides within them entirely independent and distinct from the body in which it is contained. They perceive that it is this mind, or soul, or genius, or spirit, which thinks, and desires, and decides. It commands the body, as the chief commands the slave. While the body is asleep, it is busy weaving thoughts in the sleeper’s brain, or wanders into other lands, and converses with people whom he, while awake, has never seen. They hear words of wisdom issuing from the toothless mouth of a decrepit old man. It is evident that this soul does not grow old; and therefore it does not die. The body, it is clear, is only a garment which is in time destroyed, and then where does its inmate go?

When a loved one has been taken, she haunts the memory of him who weeps till the image imprinted on the heart is reflected on the curtain of the eye. Her vision appears; not when he is quite asleep, as in an ordinary dream, but as he is passing into sleep. He meets her in the twilight land which divides the world of darkness from the world of day. He sees her form distinctly; he clasps it in his arms; he hears the accents of her sweet and gentle voice; he feels the pressure of her lips upon his own. He awakes, and the illusion is dispelled; yet with some it is so complete that they firmly believe it was a spirit whom they saw.

∗This is an extract from Chapter II, “Religion”, of The Martyrdom of Man by Winwood Reade, first published in 1872. More information on this extract can be found at http://www.hydro.ntua.gr/˜ anthony/en/martyrdom/.

Among savages it is not Love which can thus excite the imagination and deceive the sense; but reverence and fear. The great chief is dead. His vision appears in a half-waking dream: it threatens and it speaks. The dreamer believes that the form and the voice are real; and therefore he believes that the great chief still exists. It is thus that the grand idea is born. There is life after death. When the house or garment of the body is destroyed, the soul wanders forth into the air. Like the wind, it is unseen; like the wind, it can be soft and kind; like the wind, it can be terrible and cruel. The savage then believes that the pains of sickness are inflicted by the hand which so often inflicted pain upon him when it was in the flesh; and he also believes that, in battle, the departed warrior is still fighting with unseen weapons at the head of his own clan. In order to obtain the good-will of the Father-spirit, prayers are offered up to him, and food is placed beside his grave. He is, in fact, still recognised as king, and to such phantom monarchs the distinctive title of God is assigned. Each chief is deified and worshipped when he dies. The offerings and prayers are established by rule: the reigning chief becomes the family priest; he pretends to receive communications from the dead, and issues laws in their name. The deeds of valour which the chiefs performed in their lifetime are set to song; their biographies descend from generation to generation, changing in their course, and thus a regular religion and mythology are formed.

It is the nature of man to reason from himself outwards. The savage now ascribes to the various forms of matter souls or spirits, such as he imagines that he has discovered in himself. The food which he places at the grave has a soul or essence, and it is this which is eaten by the spirit of the dead, while the body of the food remains unchanged. The river is now mere water, which may dry up and perish, but there dwells within it a soul which never dies; and so with everything that lives and moves, from the blade of grass which shivers in the wind, to the star which slowly moves across the sky. But as men become more and more capable of general ideas, of classing facts into systems, and of arranging phenomena into groups, they believe in a god of the forests, a god of the waters, and a god of the sky, instead of ascribing a separate god to every tree, to every river, and to every star. Nature is placed under the dominion of a Federation of Deities. In some cases the ancestor gods are identified with these; in others, their worship is kept distinct. The trees and the animals, which were once worshipped for themselves from love or fear, are now supposed to be objects of affection to the gods, and are held sacred for their sake.

These gods are looked upon as kings. Their characters are human, and are reflected from the minds of those who have created them. Whatever the arithmetical arrangement of the gods may be—single or triune, dual or plural—they are in all countries and in all times made by man in his own image. In the plural period, some of the gods are good, and some are bad; just as there are good and evil kings. The wicked gods can be softened by flattery and presents; the good ones can be made fierce by neglect. The wicked gods obtain the largest offerings and the longest prayers, just as in despotic countries the wicked kings obtain the most liberal presents—which are merely taxes in disguise.

The savage has been led by indigestion and by dreams to believe in the existence of the soul after death, or, using simpler language, to believe in ghosts. At first these souls or ghosts have no fixed abode; they live among the graves. At a later period the savage invents a world to which the ghosts depart, and in which they reside. It is situated under ground. In that world the ghosts live precisely as they lived on earth. There is no retribution and no reward for the actions of the earthly life; that life is merely continued in another region of the world. Death is in fact regarded as a migration in which, as in all migrations, the emigrants preserve their relative positions. When a man of importance dies, his family furnish him with an outfit of slaves and wives, and pack up in his grave his arms and ornaments and clothes, that he may make his appearance in the underworld in a manner befitting his rank and fortune. It is believed that the souls of the clothes, as well as of the persons sacrificed, accompany him there; and it is sometimes believed that all the clothes which he has worn in his life will then have their resurrection day.

The under-world and the upper-world are governed by the same gods, or unseen kings. Man’s life in the upper-world is short; his life in the under-world is long. But as regards the existence of the worlds themselves, both are eternal, without beginning and without end. This idea is not a creation of the ripened intellect as is usually supposed. It is a product of limited experience, the expression of a seeming fact. The savage did not see the world begin; therefore it had no beginning. He has not seen it grow older; therefore it will have no end.

The two worlds adjoin each other, and the frontier between them is very faintly marked. The gods often dress themselves in flesh and blood and visit the earth to do evil or to do good; to make love to women, to torment their enemies, to converse with their favourites and friends. On the other hand, there are men who possess the power of leaving their bodies in their beds and of passing into the other world to obtain divine poisons which they malignantly employ. The ghosts of the dead often come and sit by their old fire-sides and eat what is set apart for them. Sometimes a departed spirit will re-enter the family, assuming a body which resembles in its features the one he previously wore. Distinguished heroes and prophets are often supposed to be hybrids or mulattoes, the result of a union between a woman and a god. Sometimes it is believed that a god has come down on earth, out of love for a certain nation, to offer himself up as a sacrifice, and so to quench the blood-thirst of some sullen and revengeful god who has that nation in his power. Sometimes a savage people believe that their kings are gods, who have deigned to take upon them a perishable body for a time; and there are countries in which a still more remarkable superstition prevails. The royal body even is immortal. The king never eats, never sleeps, and never dies. This kind of monarch is visible only to his priests. When the people wish to present a petition, he gives them audience seated behind a curtain, from beneath which he thrusts out his foot in token of assent. When he dies, he is secretly buried by the priests, and a new puppet is elected in his stead.

The savage lives in a strange world, a world of special providences and divine interpositions, not happening at long intervals and for some great end, but every day and almost at every hour. A pain, a dream, a sensation of any kind, a stroke of good or bad luck, whatever, in short, does not proceed from man, whatever we ascribe, for want of a better word, to Chance, is by him ascribed to the direct interference of the gods. He knows nothing about the laws of nature. Death itself is not a natural event. Sooner or later men make the gods angry and are killed.

It is difficult for those who have not lived among savages to perfectly realise their faith. When told that his gods do not exist, the savage merely laughs in mild wonder at such an extraordinary observation being made. It seems quite natural to him that his gods should be as his parents and grandparents have described; he believes as he breathes, without an effort; he feels that what he has been taught is true. His creed is in harmony with his intellect, and cannot be changed until his intellect is changed. If a god in a dream, or through the priests, has made him a promise and the promise is broken, he does not on that account doubt the existence of the god. He merely supposes that the god has told a lie. Nor does it seem strange to him that a god should tell a lie. His god is only a gigantic man, a sensual despotic king, who orders his subjects to give him the first fruits of the fields, the firstlings of the flock, virgins for his harem, human bodies for his cannibal repasts. As for himself, he is the slave of that god or king: he prays, that is to say, he begs; he sings hymns, that is to say, he flatters; he sacrifices, that is to say, he pays tribute, chiefly out of fear, but partly in the hope of getting something better in return—long life, riches, and fruitful wives. He is usually afraid to say of the gods what he thinks, or even to utter their real name. But sometimes he gives vent to the hatred which is burning in his heart. Writhing on a bed of sickness, he heaps curses on the god who he declares is “eating his inside”; and when converted prematurely to a higher creed, his god is still to him the invisible but human king. “Oh Allah!” a Somali woman was heard to say, “Oh Allah! May thy teeth ache like mine! Oh Allah! May thy gums be sore as mine!” That Christian monarch, the late King Peppel, once exclaimed when he thought of his approaching end, that if he could see God he would kill him at once because he made men die.

The arithmetical arrangement of the gods depends entirely upon the intellectual faculties of the people concerned. In the period of Thing-worship, as it may be termed, every brook, tree, hill and star is itself a living creature, benevolent or malignant, asleep or awake. In the next stage, every object and phenomenon is inhabited or presided over by a genius or spirit; and with some nations the virtues and the vices are also endowed with personality. As the reasoning powers of men expand, their gods diminish in number, and rule over larger areas, till finally it is perceived that there is unity in nature, that everything which exists is a part of one harmonious whole. It is then asserted that one Being manufactured the world, and rules over it supreme. But at first the Great Being is distant and indifferent; “a god sitting outside the universe;” and the old gods become viceroys to whom he has deputed the government of the world. They are afterwards degraded to the rank of messengers or angels, and it is believed that God is everywhere present; that he fills the earth and sky; that from him directly proceeds both the evil and the good. In some systems of belief, however, he is believed to be the Author of good alone, and the dominion of evil is assigned to a rebellious angel or a rival god.

So far as we have gone at present, there has been no question of morality. All doctrines relating to the creation of the world, the government of man by superior beings, and his destiny after death, are conjectures which have been given out as facts, handed down with many adornments by tradition, and accepted by posterity as “revealed religion.” They are theories more or less rational, which uncivilised men have devised, in order to explain the facts of life, and which civilised men believe that they believe. These doctrines are not in themselves of any moral value. It is of no consequence, morally speaking, whether a man believes that the world has been made by one god or by twenty. A savage is not of necessity a better man because he believes that he lives under the dominion of invisible tyrants, who will compel him, some day or other, to migrate to another land.

There is a moral sentiment in the human breast which, like intelligence, is born of obscure instincts, and which gradually becomes developed. Since the gods of men are the reflected images of men, it is evident that, as men become developed in morality, the character of their gods will also be improved. The king of a savage land punishes only offences against himself and his dependents. But when that people become more civilised, the king is regarded as the representative of public law. In the same manner the gods of a savage people demand nothing from their subjects but taxes and homage. They punish only heresy, which is equivalent to treason; blasphemy, which is equivalent to insult; and the withholding of tribute and adoration, which is equivalent to rebellion. And these are the offences which, even among civilised nations, the gods are supposed to punish most severely. But the civilised gods also require that men shall act justly to one another. They are still despots, for they order men to flatter them, and to give them money. But they are not mere selfish despots; they will reward those who do good, they will punish those who do evil to their fellow-men.

That vice should be sometimes triumphant, and virtue sometimes in distress, creates no difficulty to the savage mind. If a good man meets with misfortune, it is supposed that he is being punished for the sins of an ancestor or a relation. In a certain stage of barbarism, society is composed not of individuals, but of families. If a murder is committed, the avengers of blood kill the first man they meet belonging to the guilty clan. If the life cannot be obtained in that generation, the feud passes on, for the family never dies. It is considered just and proper that children should be punished for the sins of their fathers, unto the third and fourth generation.

In a higher state of society, this family system disappears; individualism becomes established. And as soon as this point is reached, the human mind takes a vast stride. It is discovered that the moral government of this world is defective, and it is supposed that poetical justice will be administered in the next. The doctrine of rewards and punishments in a future state comes into vogue. The world of ghosts in now divided into two compartments. One is the abode of malignant spirits, the kingdom of darkness and of pain, to which are condemned the blasphemers and the rebels, the murderers and the thieves. The other is the habitation of the gods, the kingdom of joy and light, to which angels welcome the obedient and the good. They are dressed in white robes, and adorned with golden crowns; they dwell eternally in the Royal Presence, gazing upon his lustrous countenance, and singing his praises in chorus round the throne.

To the active European mind, such a prospect is not by any means inviting; but Heaven was invented in the East; and in the East to be a courtier has always been regarded as the supreme felicity. The feelings of men towards their god, in the period at which we have now arrived, are precisely those of an Eastern subject towards his king. The Oriental king is the Lord of all the land; his subjects are his children and his slaves. The man who is doomed to death kisses the fatal firman, and submits with reverence to his fate. The man who is robbed by the king of all that he has earned will fold his hands and say, “The king gave, and the king taketh away. Blessed be the name of the king!” The man who lives in a distant province, who knows the king only by means of the taxes which are collected in his name, will snatch up his arms if he hears that his sacred person is in danger, and will defend him as he defends his children and his home. He will sacrifice his life for one whom he has never seen, and who has never done him anything but harm.

This kind of devotion is called loyalty when exhibited towards a king; piety when exhibited towards a god. But in either case the sentiment is precisely the same. It cannot be too often repeated that god is only a special name for king; that religion is a form of government, its precepts a code of laws; that priests are gatherers of divine taxes, officers of divine police; that men resort to churches to fall on their knees and to sing hymns, from the same servile propensity which makes the Oriental delight in prostrating himself before the throne; that the noble enthusiasm which inspires men to devote themselves to the service of their god, and to suffer death rather than deny his name, is identical with the devotion of the faithful subject who, to serve his royal master, gives up his fortune or his life without the faintest prospect of reward. The religious sentiment, about which so much has been said, has nothing distinctive in itself. Love and fear, self-denial and devotion, existed before those phantoms were created which men call gods; and men have merely applied to invisible kings the sentiments which they had previously felt towards their earthly kings. If they are a people in a savage state, they hate both kings and gods within their hearts, and obey them only out of fear. If they are a people in a higher state, love is mingled with their fear, producing an affectionate awe which, in itself, is pleasing to the mind. That the worship of the unseen king should survive the worship of the earthly king is natural enough; but even that will not endure for ever; the time is coming when the crowned idea will be cast aside and the despotic shadow disappear.

By thus translating, or by retranslating, god into king, piety into loyalty, and so on; by bearing in mind that the gods were not abstract ideas to our ancestors as they are to us, but bonâ-fide men, differing only from men on earth in their invisibility and other magic powers; by noting that the moral disposition of a god is an image of the moral sense of those who worship him— their beau-ideal of what a king should be; that the number and arrangement of the gods depend exclusively on the intellectual faculties of the people concerned, on their knowledge of nature, and perhaps, to some extent, on the political forms of government under which they live; above all by remembering that there is a gradual development in supernatural ideas, the student of comparative religion will be able to sift and classify with ease and clearness dense masses of mythology. But he must understand that the various stages overlap. Just as sailing vessels and four-horse coaches are still used in this age of steam, and as stone implements were still to be found in use long after the age of iron had set in, so in the early period of god belief, thing-worship still to a certain extent endured. In a treaty between Hannibal and Philip of Macedonia, which Polybius preserved, the contracting parties take oath with one another “in the presence of Jupiter, Juno, and Apollo; in the presence of the Deity of the Carthaginians and of Hercules and of Iolaus; in the presence of Mars, Triton, and Neptune; in the presence of all the gods who are with us in the camp; and of the sun, the moon, and the earth; the rivers, the lakes, and the waters.” In the time of Socrates the Athenians regarded the sun as an individual. Alexander, according to Arrian, sacrificed not only to the gods of the sea, but “the sea itself was honoured with his munificence.” Even in Job, the purest of all monotheistic works, the stars are supposed to be live creatures which sing around the heavenly throne.

Again, in those countries where two distinct classes of men exist, the one intellectual and learned, the other illiterate and degraded, there will be in reality two religions though nominally there may be only one. Among the ancient Sabæans the one class adored spirits who inhabited the stars, the other class adored the stars themselves. Among the worshippers of fire, that element to one class was merely an emblem, to the other an actual person. Wherever idols or images are used, the same phenomenon occurs. These idols are intended by the priests as aids to devotion, as books for those who cannot read. But the savage believes that his god inhabits the image, or even regards the image as itself a god. His feelings towards it are those of a child towards her doll. She knows that it is filled with saw-dust and made of painted wood, and yet she loves it as if it were alive. Such is precisely the illusion of the savage, for he possesses the imagination of a child. He talks to his idol fondly and washes its face with oil or rum; beats it if it will not give him what he asks; and hides it in his waist-cloth if he is going to do something which he does not wish it to see.

There is one other point which it is necessary to observe. A god’s moral disposition, his ideas of right and wrong, are those of the people by whom he is created. Wandering tribes do not, as a rule, consider it wrong to rob outside the circle of their clan; their god is therefore a robber like themselves. If they settle in a fertile country, pass into the agricultural state, build towns, and become peaceful citizens with property of their own, they change their views respecting theft, and accordingly their god forbids it in his laws. But it sometimes happens that the sayings and doings of the tent-god are preserved in writings which are accepted as revelation by the people of a later and a better age. Then may be observed the curious and by no means pleasing spectacle of a people outgrowing their religion, and believing that heir god performed actions which would be punished with the gallows if they were done by men.

The mind of an ordinary man is in so imperfect a condition that it requires a creed; that is to say, a theory concerning the unknown and the unknowable in which it may place its deluded faith and be at rest. But whatever the creed may be, it should be one which is on a level with the intellect, and which inquiry will strengthen, not destroy.

As for minds of the highest order, they must ever remain in suspension of judgment and in doubt. Not only do they reject the absurd traditions of the Jews, but also the most ingenious attempts which have been made to explain, on rational and moral grounds, the origin and purpose of the universe. Intense and longcontinued labour reveals to them this alone: that there are regions of thought so subtle and so sublime that the human mind is unable therein to expand its wings, to exercise its strength. But there is a wide speculative field in which man is permitted to toil with the hope of rich reward, in which observation and experience can supply materials to his imagination and his reason. In this field two great discoveries have been already made. First, that there is a unity of plan in nature; that the universe resembles a body in which all the limbs and organs are connected with one another; and secondly, that all phenomena, physical and moral, are subject to laws as invariable as those which regulate the rising and setting of the sun. It is in reality as foolish to pray for rain or a fair wind as it would be to pray that the sun should set in the middle of the day. It is as foolish to pray for the healing of a disease or for daily bread as it is to pray for rain or a fair wind. It is as foolish to pray for a pure heart or for mental repose as it is to pray for help in sickness or misfortune. All the events which occur upon the earth result from Law; even those actions which are entirely dependent on the caprices of the memory, or the impulse of the passions, are shown by statistics to be, when taken in the gross, entirely independent of the human will. As a single atom, man is an enigma; as a whole, he is a mathematical problem. As an individual, he is a free agent; as a species, the offspring of necessity.

The unity of the universe is a scientific fact. To assert that it is the operation of a single Mind is a conjecture based upon analogy, and analogy may be a deceptive guide. It is the most reasonable guess that can be made, but still it is no more than a guess; and it is one by which nothing after all is really gained. It tells us that the earth rests upon the tortoise; it does not tell us on what the tortoise rests. God issued the laws which manufactured the universe, and which rule it in his growth. But who made God? Theologians declare that he made himself; and materialists declare that Matter made itself; and both utter barren phrases, idle words. The whole subject is beyond the powers of the human intellect in its present state. All that we can ascertain is this: that we are governed by physical laws which it is our duty as scholars of Nature to investigate; and by moral laws which it is our duty as citizens of Nature to obey.